OpenStudy (anonymous):

If a plane can travel 470mph with wind and 430mph against the wind, find the speed of the plane in still air. What is the speed of the plane is in still air?

OpenStudy (lgbasallote):

downstream = speed of plane + speed of wind upstream = speed of plane - speed of wind knowing these things... let x = speed of plane y = speed of wind 470 mph = x + y 430 mph = x - y using systems of linear equations... 900 = 2x x = 450 mph is that the answer you're looking for?

OpenStudy (anonymous):

That's I thought. But wasn't sure. Thank u